A 450$ stereo after a 15% discount
Found 2 solutions by lynnlo, ikleyn:
Answer by lynnlo(4176)
Answer by ikleyn(53765)

.
A 450$ stereo after a 15% discount
~~~~~~~~~~~~~~~~~~~~~~~~

It is 450 - 0.15*450 = 450*(1-0.15) = 0.85*450 = 382.50 dollars.    ANSWER

Solved.